Suppose you observe the following situation:

Security Beta Expected Return
Peat Co. 1.15 10.0
Re-Peat Co. 0.90 9.0

Assume these securities are correctly priced. Based on the CAPM, what is the expected return on the market? What is the risk-free rate? (Do not round intermediate calculations. Enter your answers as a percent rounded to 2 decimal places.)

Expected return on market %
Risk-free rate %
Answer #1

Expected Return=risk-free rate +Beta*(market rate- risk-free rate)

10=Rf+1.15*(Rm-Rf)

10=1.15Rm-0.15Rf

Rm=(10+0.15Rf)/1.15

Also:

9=Rf+0.9*(Rm-Rf)

9=0.9Rm+0.1Rf

9=0.9*(10+0.15Rf)/1.15+0.1Rf

9=7.826086957+0.117391304 Rf+0.1 Rf

Rf=(9-7.826086957)/(0.117391304+0.1)

=5.4%(Approx)=risk free rate

Rm=(10+0.15Rf)/1.15

=9.4%(Approx)=market rate